Edited: 02/11/14, 11:49
Minutes of the February 11, 2014 weekly meeting
Present: A. Camsonne, C. Desnault, N. Israel, M. Dlamini, K. Park, Frederic Georges. , M. Carmignotto, J. Roche
- Calorimeter in the test lab: report by Kijun
The work was performed by Hashir, Lee and Kiun with some help from Julie.
- Since the last meeting, all blocks of the calo have been tested using cosmic rays. The blocks were powered one by one. All but 5 blocks/PMT/ampli were found to be operational. For the 5 "bad" ones, the HV and signal cables were interchanged with neighboring blocks (up to the cables between the distribution boards and the bases). The problems are with the pmt and/or ampli. These 5 blocks will have to be pulled out of the assembly for further investigations.
- Preparing to work on the paddle for the cosmic paddle. The screws that stuck out in the space where the cosmics paddle will go, have been replaced.
- This week will be spend establishing the trigger, cabling the ARS and setting up the DAQ.
- Trigger: report by Norman
- Norman confirmed that opening the front panel in order to cool the modules did not make the module work. (i.e. setting the front panel back did not break the module)
- Coda has been installed on the laptop Norman uses in the electronic lab. It is almost functional but for using the correct VME list that allow the full reading of the trigger header.
- Plan for next week:
- Implement run with block transfer
- Investigate the source of the zeros in the ADC readout. Norman will put together a list of tests he plans before performing them.
- Norman has some questions out to Magali (regarding some LED being red and the zeros in the read out). Hoping for answers very soon.
- LED : report by Marco
- will be at JLab next week
- got some feedback from Magali about the LED operation. As a result went on building a lemo to rub ban cable necessary
Please send comment and correction to Julie
|